OpenStudy (anonymous):

You hear sounds at the following decibel levels: 5 dB, 10 dB, 20 dB, and 40 dB. Which sound is 10 times higher than the lowest pressure humans can hear? A. 5 dB B. 10 dB C. 20 dB D. 40 dB **Not sure how to find this!! :( Thank you!!:)

OpenStudy (michele_laino):

we have to use this definition: \[\Large N = 10{\log _{10}}\left( {\frac{{{P_1}}}{{{P_0}}}} \right)\] where N is the number of decibels, and P_0 is a reference pressure

OpenStudy (anonymous):

oh okay! how can we find what to plug in?

OpenStudy (michele_laino):

we have to know how much is the lowest pressure

OpenStudy (anonymous):

ohhh would the lowest pressure be the 5 dB?

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

or wait, it is 5 dB? :/

OpenStudy (michele_laino):

no, sorry, we can answer to your problem without know how much is the lowest pressure. Since we can write the lowest pressure as our reference pressure, and we can set P_1 = 10*P_0

OpenStudy (anonymous):

ohh okay! :)

OpenStudy (michele_laino):

so we have: \[\Large \begin{gathered} {P_1} = 10{P_0} \hfill \\ N = 10{\log _{10}}\left( {\frac{{10{P_0}}}{{{P_0}}}} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

10?

OpenStudy (michele_laino):

ok! 10 dB

OpenStudy (anonymous):

ohh so that is our solution? 10 dB? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

woo! thank you!!:D

OpenStudy (michele_laino):

thanks! :)